OpenStudy (anonymous):

10x201 divided by 20

OpenStudy (anonymous):

Is this Order of operations?

OpenStudy (anonymous):

10 x 201 = 2010 2010 / 20 = ?

OpenStudy (anonymous):

do the multiplying first, then the division

OpenStudy (anonymous):

10.5 is the answer 10*201=210 210/20 - 10.5

OpenStudy (anonymous):

no wait

OpenStudy (anonymous):

its 100.5 my bad

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

10*201= 2,010 2,010 / 20 =100.5

OpenStudy (anonymous):

ty